<<TableOfContents(3)>>

= Agenda/Minutes =
[[https://zoom.us/j/203743987|Zoom Meeting]]

----------
= CRY+VIS Interface Meeting =
Date: 2021/10/15 13:00-13:25

Participants: Takahashi, Hirata, Miyoki, Yasui, Ikeda, Ushiba, Uchiyama, Yokozawa, Sato, Node

[[http://gwwiki.icrr.u-tokyo.ac.jp/JGWwiki/KAGRA/Upgrades/PostO3Upgrades|Post O3 Upgrade portal]]

== Schedule ==
 * IX: Health check. F0 behavire.
 * EX: Health check.
 * IY: (LVDT driver. Cable check.)
 * EY: H4 geophone. F0 FR. Height adjustment. LVDT driver. IP position. Insulator repair.

== Status ==
 * TM height setting: ITMY
 * Commissioning in air: Health check.
 * Commissioning in vacuum:
 * Overall commissioning:
 * Acceptance check:
 * PF actuator: At the in-vacuum commissioning.
 * New vacuum pumps for V-chambers: Under the bidding process.
 * Heater control: To PRs in the next.

== Discussion ==
----------
= VIS Meeting =
Date: 2021/10/15 14:00-14:40

Participants: Takahashi, Ushiba, Ishizaki, Hirata, Uchiyama, Sugiyama, Terrence, Fabian, Miyoki, Tanaka, Sato

== Summary Report ==
=== Type-A ===
 * EY: Installation of the H4 geophone. Modification of the F0 FR. Height adjustment with the jacks. Replacement of the LVDT driver for IP. Adjustment of the IP position.
 * IX: Repair of the cable connection in F0.

=== Type-B ===
 * BS: We tested the robustness of the outer frame
  * Resolution of the dial gauge: 0.01 mm.
  * Weight added to the frame: ~71 kg.
  * '''Corner (-X,-Y) : -3 um '''
  * '''Corner (-X,+Y): -10um'''
  * '''Corner (+X,+Y): it didn't seem to move much.'''
  * '''Corner (+X,-Y): -8 um'''
 * SR2: The geophone theoretical noise was calculated (shown in this meeting last year also). '''The noise is higher than expected.'''
 * SR3: The geophone theoretical noise was calculated.
  * '''Below 500 mHz, the calculated and measured values seem to coincide. '''
  * As  far as I remember being told, these geophones have been very noisy   after installation on the IP table. I need to measure the IP free swing   displacement ASD on a day with low micro-seismic in order to compare.

=== Type-Bp ===
 * SR2: We fixed the BF picomotor cable.
 * PR3:
  * On  monday we moved the V1 OSEM within range. We left it at ~9940 cnt.  Although I don't expect this to happen, let's see if the OSEM remains in  range after the next eathquake.
  * We assesed whether is possible to lower the SF keystone by 5.2 mm: it seems possible.
   * The  concern was cables underneath the BF touching the security structure  ring, but there are functional gaps after we lowered it.
   * First we moved the keystone using the lock screws, but later we used 405 grams.
   * Lowering the SF yoke-coil assembly might damage cables around, but I guess we can fix them if such a thing happens.
   * It's better to add mass on the keystone rather than on the BF. Hirata-san has designed an adapter.

=== Type-C ===
 * Unification of the coil name notation (H1, H2, ..) was suggested.

=== Electronics ===
=== Control ===
== Schedule ==
See [[http://gwwiki.icrr.u-tokyo.ac.jp/JGWwiki/KAGRA/Subgroups/VIS/Schedule|Schedule]]

 * Type-A team: From the satellite room in Mitaka.
 * Type-B team: Geophone noise, IP LVDT.

== Safety ==
== Discussion ==
== Travel Plans ==
(Date in Kamioka)

The week from 10/18

 * Hirata: 10/19-22
 * Yano: 10/18-22

The week from 10/25

 * Takahashi: 10/26-29
 * Sato: 10/26-29
 * Hirata: 10/26-29
 * Yano: 10/25-29

== Next meeting ==
 * 2021/10/22(FRI)